# Rolling in the Shadows: Analyzing the Extraction of MEV Across Layer-2 Rollups

### v1

@@ Abstract.
Ethereum, Layer 2, Rollups, MEV, Cross-Layer Sandwiching
@@ 1. Introduction
Maximal Extractable Value (MEV) has become an essential part of Ethereum [56, 53]. The idea behind MEV is to extract value (i.e., monetary profit) from transactions performed by other entities on the blockchain by influencing their order of execution. Several strategies have been proposed, the most popular ones being arbitrage, liquidation, and sandwiching [56, 53]. MEV is a double-edged sword. While strategies such as arbitrage and liquidation help maintaining markets healthy across different DeFi protocols (e.g., liquidation of unhealthy loans, price balancing across exchanges, etc.), they can also have negative side effects. For example, they result in increased transaction fees as a result of gas price auctions [20] or wasteful transactions due to failed attempts to extract MEV. Some strategies, such as sandwiching, are even considered entierly destructive. Sandwiching occurs, for example, by placing a buy order right before a pending trade (i.e., frontrunning) and a sell order right after it (i.e., backrunning), thereby forming a “sandwich” and resulting in traders making less profit on their trades.
@@
In this paper, we shed light into how much MEV is being extracted on popular rollups such as Arbitrum [7], Optimism [44], and zkSync [67], and how this compares to Ethereum in terms of volume, profit, and costs across a 32 month period. We also analyze the usage of flash loans, code reuse and competition among MEV extractors as well as the time extractors take to respond to MEV opportunities. Furthermore, we check for traditional sandwiching on rollups, confirming that so far sequencers behave as intended and do not perform any evident sandwiching. Nonetheless, after carefully studying the way transactions are sent across rollups and Ethereum, we propose and evaluate three novel cross-layer sandwich attacks that normal users can leverage to mount sandwich attacks on rollups. These attacks exploit transactions that perform trades on rollups but which are emitted via Ethereum. Our simulation using past mainnet data reveals that attackers could have made roughly 2 million USD profit. Finally, we also discuss the generalization of our attacks and potential countermeasures.
Ethical Considerations. For ethical reasons, we demonstrate the feasibility of our attacks by executing them only on the testnet, thereby targeting only our own victim transactions. On the mainnet, we only perform a local simulation of our attacks and do not actually carry out out the attacks by broadcasting the transactions.
@@ 3.3. Detecting Sandwiching
For each detected transfer event we aim to find its counterpart, namely a transfer event that was emitted afterwards via a different transaction by the same token smart contract but which transfers the tokens in the opposite direction. Hence, we search for transactions $T_{A_{1}}$ and $T_{A_{2}}$, such that $T_{A_{2}}$ has a larger transaction index value than $T_{A_{1}}$ and where the sender of $T_{A_{1}}$ is equivalent to the receiver of $T_{A_{2}}$ and the receiver of $T_{A_{1}}$ is equivalent to the sender of $T_{A_{2}}$. Moreover, the amount transferred in $T_{A_{2}}$ has to be either the same or smaller than the amount transferred in $T_{A_{1}}$. The same applies for rollups, except that we verify the order by first checking batch height and then within the same batch the transaction index. Once such a pair has been identified, we search for one or multiple victim transactions (i.e., $T_{V}$), by checking if there are any token transfer events emitted by the same token in between the identified attacker events, which have the same sender as $T_{A_{1}}$ but a different receiver .
@@ 3.7. Limitations
Our methodology currently does not capture all MEV as it searches for hard-coded events triggered by a specific set of DeFi protocols. Our results should be considered a lower bound on the volume of MEV extraction as well as a comparison of the MEV extraction between several popular DeFi protocols. Similar to prior works [53, 40, 56, 61], our price calculation relies on CoinGecko’s public API [14] to identify the value of a token at a given time. As a result, the profit calculation might not precisely reflect the actual profit that the extractor made on that day. Moreover, CoinGecko only tracks popular tokens, thus, for less popular tokens we are unable to calculate the profits made by the extractor.
@@ 4.1. Volume
For each of the rollups, we see gradual but significant increases in the number of liquidations and arbitrages over the course of our measurement period. In the bottom plot of Figure 5 we see a lot of variability in the number of liquidations on Ethereum, vacillating between 498 and 13,121 for each four month period. For Arbitrum and Optimism, we see rapid increase in the early months after the technology was deployed. Arbitrum then stabilizes between 309 and 665 liquidations per month. Optimism appears to decrease after its peak of 1,307 liquidations in the summer of 2022. Interestingly, there is a decline in the number of liquidations on Ethereum between June 2022 and June 2023. This could be a result of a smaller number of loan or broader economic trends causing the collaterals to maintain their value with higher probability.
We see similar trends in the top plot of Figure 5. The salient difference between the two plots, though, is that all three of the rollups eventually surpass Ethereum in the number of detected arbitrages. In the month of April 2023, Arbitrum had $7.2$$\times$ more arbitrages than Ethereum, although the difference narrows in the subsequent months. Again, this is likely related to broader economic trends and the increasing usage of rollups as the transaction fees tend to be much lower. After opening up to developers in February 2023, the first MEV was detected around March 25, 2023 on zkSync, highlighting the rapid adoption of MEV extractors even on ZK-based rollups. Finally, we found no instances of sandwiching on any of the rollup platforms suggesting that the sequencers are not performing sandwich trades. We provide additional granularity in Appendix C. However, in Section 5, we show that cross-layer sandwiching is a viable threat vector to traders on rollups. Cross-layer sandwich attacks can be mounted by normal users and do not require control over pending transactions.
@@ 4.2. Profits
Strategy Profit Ethereum Arbitrum Optimism zkSync Arbitrage Total $209981692.18$ $17419113.84$ $2514046.36$ $640243.36$ Max $25614599.91$ $1913418.44$ $79927.65$ $23891.42$ $\text{P}_{90}$ $87.51$ $6.55$ $1.94$ $7.52$ Mean $74.12$ $10.14$ $2.21$ $6.12$ Median $4.80$ $0.27$ $0.23$ $0.83$ Min $-15828.08$ $-1659.02$ $-15963.71$ $-21.39$ Liquidation Total $232498892.47$ $856503.70$ $1308358.38$ $10733.13$ Max $2675295.15$ $71224.66$ $288573.28$ $4014.54$ $\text{P}_{90}$ $4814.00$ $226.37$ $74.84$ $59.87$ Mean $5323.75$ $206.44$ $296.08$ $42.59$ Median $116.91$ $1.60$ $0.67$ $4.98$ Min $-127586.84$ $-821.46$ $-2422.01$ $-1682.03$ Total $442480584.65$ $18275617.54$ $3822404.74$ $650976.49$
@@
In Table 3, we see that Ethereum is still the platform with the highest profits, both cumulatively and on average. In fact, the cumulative profits from MEV on Ethereum are close to 10$\times$ the profits on the other three platforms combined for both arbitrage and liquidation. This is likely because Ethereum is simply more popular and thus has more pools and more opportunities and more liquidity. Additionally, we note that the mean values are much higher than the median values while the maximum values are much higher than the P90 values for all platforms—this indicates that a small percentage of transactions account for a disproportionate fraction of the total profits. So while by definition, 90% of transactions are below the P90 in terms of profits, those that are above the 90th percentile are way above. Among the rollups, Aribtrum has significantly greater arbitrage profits than either Optimism or zkSync. While for liquidations, Optimism has slightly higher profits. We suspect that this is due to the availability of different contracts and services (including loans) between these platforms.
@@ 4.4. Opportunities
Figure 7 depicts the cumulative distribution of the block distance between extracted MEV and MEV opportunities for up to a distance of 100 blocks. We see that for Ethereum, over 50% of the MEV extractions occurred within the same block as the opportunity. Interestingly, for zkSync and former releases of Arbitrum and Optimism (i.e., Classic and Pre-Bedrock, respectively) we find that the smallest block distance is one (i.e., the MEV extraction never occurred in the next block as the opportunity). However, for newer releases of Arbitrum and Optimism (i.e., Nitro and Post-Bedrock, respectively) we do find a small amount of MEV extractions that occur within the same block. This is an interesting result since newer releases of Arbitrum and Optimism still do not have a public mempool. Yet, some extractors were able to insert their transactions within the same block as the opportunity. For instance, according to our dataset, address 0x4b9406…5e3021 performed a total of six successful liquidations on Optimism, where all liquidations were performed within the same block as the opportunity, which hints towards a systematic approach. One example is block $108560468$, which contains 18 transactions and the extractor’s transaction was included at position 11 whereas the opportunity is included at position 4. The block was built on August 22, 2023 i.e., towards the end of our measurement period.
@@ 4.7. Code Reuse
Table 5. MEV bots with identical bytecode across chains.
@@ 5.3.1. Victim Inference
Once linked, we scan the L2 transactions for DEX swaps. A swap typically results in two ERC-20 “Transfer” events being emitted. Hence, we scan the events emitted by the identified L2 transactions and mark these as potential victims if we find two “Transfer” events emitted by the same token smart contract, where the sender and receiver of the tokens are swapped. We found a total of $1916524$ L2 transactions emitted via L1, out of which $1459570$ were observed on Arbitrum, $424212$ on Optimism, and $32742$ on zkSync. From the observed 1.9M L2 transactions, only $170674$ ($\approx$ 9%) can be considered potential victims based on our methodology to detect token swaps, where $87921$ are from Arbitrum and $82753$ from Optimism. We could not find potential victims on zkSync. This is most likely due to the fact that there was no cross-layer DEX currently deployed on zkSync at the time of our measurement. Appendix B provides a detailed overview of the identified DEXes and their liquidity pools on Arbitrum and Optimism. We observe that essentially all potential victim transactions originate from users using the Hop Protocol [51]. The Hop protocol allows users to send tokens from one rollup or chain to another by employing their own market makers, hence making it a perfect victim for our cross-layer sandwich attacks.
@@ 5.3.2. Transaction Inclusion Delays
Appendix D provides a detailed overview on the profits obtained for each attack strategy across Arbitrum and Optimism using different attacker budgets. Interestingly, not all identified potential victim transactions are profitable. These transactions are either too small to offer revenue outweighing costs, or have low slippage tolerance settings which do not provide enough room for frontrunning. In Table 6, we observe that the number of profitable transactions increases together with the attacker’s capital, suggesting that several victim transactions require large purchases by attackers in order to be profitable. Nonetheless, we can state that the maximum total profit is 1.2M USD for Arbitrum and 767K USD for Optimism. Figure 10 further highlights that Optimism is more profitable than Arbitrum for attackers that posses a capital up to 1M USD. In addition, we observe that attack strategy $S_{3}$ is always more profitable than $S_{2}$ and $S_{1}$. However, the difference diminishes as the attacker capital increases. Hence, we conclude that either strategies are profitable, but that for attackers with smaller capital (e.g., 1K USD) strategy $S_{3}$ is more profitable and relevant than for attackers with larger capital (e.g., 1M USD).
@@ 5.4. Testnet Validation
While our mainnet simulation provides insights such as number of potential victims and estimated profits, it does not validate the deployability of our attacks in practice. To that end, we leverage live test networks to demonstrate the practicality of our cross-layer sandwich attacks. Given the similarities between testnets and mainnets, we can validate the feasibility of our attacks in an ethical manner without impacting real users and requiring substantial financial resources. For our analysis, we used the Sepolia testnet alongside its corresponding L2 counterparts for Arbitrum, Optimism, and zkSync. We developed a script to automate the process of deploying our own attacker and victim accounts on each chain including our own Uniswap V2-inspired DEXes on Arbitrum, Optimism, and zkSync. The only victim was our own account even on the testnet. The script automatically performs the three attack strategies. First, a victim transaction $T_{V}$ is sent via the official L1 bridge of the rollup that is under test. The victim transaction simply performs a swap on the Uniswap-alike DEX that the script previously deployed on L2. Afterwards, the script simulates the attacker transactions $T_{A_{1}}$ and $T_{A_{2}}$ according to the selected strategy. The script either scans the mempool or the latest blocks on L1 and searches for our victim transaction $T_{V}$ as detailed in Section 5.3.1 and crafts the corresponding transactions $T_{A_{1}}$ and $T_{A_{2}}$, which essentially buy and sell the same asset as $T_{V}$ on the Uniswap-alike DEX on L2. See Appendix E for a detailed list of our successfully deployed attack transactions across Arbitrum, Optimism, and zkSync.
@@ 6.1. Generalization and Limitations
All of our proposed cross-layer sandwich attacks exploit the fact that users submit L2 transactions via L1. While this seems counter-intuitive, there exist legitimate use cases which require such behavior. For example, they are used in cross-chain swaps as provided via the Hop Protocol [51] or the Across Protocol [50]. The reason why we were only able to find victim transactions related to Hop, was due to the fact that Across leverages their own bridges and does not rely on the standard bridges deployed by individual rollup technologies. However, Across’s bridge also hinges on emitting events to trigger cross-chain swaps, hence, our victim inference could be adjusted for Across as well. Attacks $S_{1}$ and $S_{2}$ depend on the fact that attackers can observe victim transactions via the mempool. Hence, any L1 blockchain that has a public mempool is susceptible to these two strategies. Finally, we did not find any victim transactions on zkSync due to the fact that during our data collection period no cross-chain exchange such as Hop was operating.

### v2

@@ Abstract.
Ethereum; layer-2; rollups; MEV; cross-layer; sandwiching; DeFi
@@ 1. Introduction
Maximal Extractable Value (MEV) has become an essential part of Ethereum [60, 57]. The idea behind MEV is to extract value (i.e., monetary profit) from transactions performed by other entities on the blockchain by influencing their order of execution. Several strategies have been proposed, the most popular ones being arbitrage, liquidation, and sandwiching [60, 57]. MEV is a double-edged sword. While strategies such as arbitrage and liquidation help maintaining markets healthy across different DeFi protocols (e.g., liquidation of unhealthy loans, price balancing across exchanges, etc.), they can also have negative side effects. For example, they result in increased transaction fees as a result of gas price auctions [22] or wasteful transactions due to failed attempts to extract MEV. Some strategies, such as sandwiching, are even considered entirely destructive. Sandwiching occurs, for example, by placing a buy order right before a pending trade (i.e., frontrunning) and a sell order right after it (i.e., backrunning), thereby forming a “sandwich” and resulting in traders making less profit on their trades.
@@
In this paper, we shed light into how much MEV is being extracted on popular rollups such as Arbitrum [7], Optimism [47], and zkSync [74] as these are the rollups with the highest market share in terms of total value locked at the time of writing [39]. We analyze how MEV extraction on these rollups compares to Ethereum in terms of volume, profit, and costs across a 32 month period. We also analyze the usage of flash loans, code reuse and competition among MEV extractors as well as the time extractors take to respond to MEV opportunities. Furthermore, we check for traditional sandwiching on rollups, confirming that so far sequencers behave as intended and do not perform any evident sandwiching. Nonetheless, after carefully studying the way transactions are sent across rollups and Ethereum, we propose and evaluate three novel cross-layer sandwich attacks that normal users can leverage to mount sandwich attacks on rollups. These attacks exploit transactions that perform trades on rollups but which are emitted via Ethereum. Our simulation using past mainnet data reveals that attackers could have made roughly 2 million USD profit. Finally, we also discuss the generalization of our attacks and potential countermeasures.
Ethical Considerations. For ethical reasons, we demonstrate the feasibility of our attacks by executing them only on the testnet, thereby targeting only our own victim transactions. On the mainnet, we only perform a local simulation of our attacks and do not actually carry out the attacks by broadcasting the transactions. The attacks performed on the testnet did not rely on real victim transactions, but on victim transactions emitted by us. Our simulation on real victim transactions on the mainnet cannot have an impact anymore on these victims as their transactions are already finalized on L1 and L2, and cannot be exploited anymore.
@@ 3.3. Detecting Sandwiching
For each detected transfer event we aim to find its counterpart, namely a transfer event that was emitted afterwards via a different transaction by the same token smart contract but which transfers the tokens in the opposite direction. Hence, we search for transactions $T_{A_{1}}$ and $T_{A_{2}}$, such that $T_{A_{2}}$ has a larger transaction index value than $T_{A_{1}}$ and where the sender of $T_{A_{1}}$ is equivalent to the receiver of $T_{A_{2}}$ and the receiver of $T_{A_{1}}$ is equivalent to the sender of $T_{A_{2}}$. Moreover, the amount transferred in $T_{A_{2}}$ has to be either the same or smaller than the amount transferred in $T_{A_{1}}$. The same applies for rollups, except that we verify the order by first checking batch height and then within the same batch the transaction index. Once such a pair has been identified, we search for one or multiple victim transactions (i.e., $T_{V}$), by checking if there are any token transfer events emitted by the same token in between the identified attacker events, which have the same sender as $T_{A_{1}}$ but a different receiver. The sender for $T_{A_{1}}$ and $T_{V}$ are expected the same since it is the address of the same DEX, however the receiver must be different since it is the address of the actual attacker/victim.
@@ 3.7. Limitations
Our methodology currently does not capture all MEV as it searches for hard-coded events triggered by a specific set of DeFi protocols, thus we will miss MEV emitted by other protocols. Hence, our results should be considered as a lower bound on the volume of extraction as well as in terms of comparison between DeFi protocols. Moreover, our competition measurement does not capture MEV transactions that did not make it on-chain. Similar to prior works [57, 43, 60, 66], which measured MEV on L1, our price calculation relies on CoinGecko’s public API [16] to identify the value of a token at a given time. As a result, the profit calculation might not precisely reflect the actual profit that the extractor made on that day. Moreover, CoinGecko only tracks popular tokens, thus, for less popular tokens we are unable to calculate the profits made by the extractor.
@@ 4.1. Volume
For each of the rollups, we see gradual but significant increases in the number of liquidations and arbitrages over the course of our measurement period. In the bottom plot of Figure 5 we see a lot of variability in the number of liquidations on Ethereum, vacillating between 498 and 13,121 for each four month period. For Arbitrum and Optimism, we see rapid increase in the early months after the technology was deployed. Arbitrum then stabilizes between 309 and 665 liquidations per month. Optimism appears to decrease after its peak of 1,307 liquidations in the summer of 2022. Interestingly, there is a decline in the number of liquidations on Ethereum between June 2022 and June 2023. This could be a result of a smaller number of loans or broader economic trends causing the collaterals to maintain their value with higher probability.
We see similar trends in the top plot of Figure 5. The salient difference between the two plots, though, is that all three of the rollups eventually surpass Ethereum in the number of detected arbitrages. In the month of April 2023, Arbitrum had $7.2$$\times$ more arbitrages than Ethereum, although the difference narrows in the subsequent months. Again, this is likely related to broader economic trends and the increasing usage of rollups as the transaction fees tend to be much lower. After opening up to developers in February 2023, the first MEV was detected around March 25, 2023 on zkSync, highlighting the rapid adoption of MEV extractors even on ZK-based rollups. Finally, we found no instances of sandwiching on any of the rollup platforms suggesting that the sequencers are not performing sandwich trades. We provide additional granularity in Table 8 in Appendix B. However, in Section 5, we show that cross-layer sandwiching is a viable threat vector to traders on rollups. Cross-layer sandwich attacks can be mounted by normal users and do not require control over pending transactions.
@@ 4.2. Profits
Strategy Profit Ethereum Arbitrum Optimism zkSync Arbitrage Total $209981692.18$ $17419113.84$ $2514046.36$ $640243.36$ Max $25614599.91$ $1913418.44$ $79927.65$ $23891.42$ $\text{P}_{90}$ $87.51$ $6.55$ $1.94$ $7.52$ Mean $74.12$ $10.14$ $2.21$ $6.12$ Median $4.80$ $0.27$ $0.23$ $0.83$ Min $-15828.08$ $-1659.02$ $-15963.71$ $-21.39$ Liquidation Total $232498892.47$ $856503.70$ $1308358.38$ $10733.13$ Max $2675295.15$ $71224.66$ $288573.28$ $4014.54$ $\text{P}_{90}$ $4814.07$ $225.64$ $74.82$ $60.10$ Mean $5323.75$ $206.44$ $296.08$ $42.59$ Median $116.91$ $1.60$ $0.67$ $4.98$ Min $-127586.84$ $-821.46$ $-2422.01$ $-1682.03$ Total $442480584.65$ $18275617.54$ $3822404.73$ $650976.49$
@@
In Table 3, we see that Ethereum is still the platform with the highest profits, both cumulatively and on average. In fact, the cumulative profits from MEV on Ethereum are close to 10$\times$ the profits on the other three platforms combined for both arbitrage and liquidation. This is likely because Ethereum is simply more popular and thus has more pools and more opportunities and more liquidity. Additionally, we note that the mean values are much higher than the median values while the maximum values are much higher than the P90 values for all platforms—this indicates that a small percentage of transactions account for a disproportionate fraction of the total profits. So while by definition, 90% of transactions are below the P90 in terms of profits, those that are above the 90th percentile are way above. Among the rollups, Aribtrum has significantly greater arbitrage profits than either Optimism or zkSync. While for liquidations, Optimism has slightly higher profits. We suspect that the arbitrage profits on Arbitrum are higher due to users performing trades with higher amounts as compared to Optimism. We also suspect that liquidation profits are higher on Optimism due to higher collaterals.
@@ 4.4. Opportunities
Figure 7 depicts the cumulative distribution of the block distance between extracted MEV and MEV opportunities for up to a distance of 100 blocks. We see that for Ethereum, over 50% of the MEV extractions occurred within the same block as the opportunity (i.e., block distance of 00). Interestingly, for zkSync and former releases of Arbitrum and Optimism (i.e., Classic and Pre-Bedrock, respectively) we find that the smallest block distance is 1 (i.e., the MEV extraction never occurred in the same block as the opportunity). However, for newer releases of Arbitrum and Optimism (i.e., Nitro and Post-Bedrock, respectively) we do find a small amount of MEV extractions that occur within the same block. This is an interesting result since newer releases of Arbitrum and Optimism still do not have a public mempool. Yet, some extractors were able to insert their transactions within the same block as the opportunity. For instance, according to our dataset, address 0x4b9406…5e3021 performed a total of six successful liquidations on Optimism, where all liquidations were performed within the same block as the opportunity, which hints towards a systematic approach. One example is block $108560468$, which contains 18 transactions and the extractor’s transaction was included at position 11 whereas the opportunity is included at position 4. The block was built on August 22, 2023 i.e., towards the end of our measurement period.
@@ 4.7. Code Reuse
Table 5. MEV bots with identical bytecode across chains.
@@ 5.3.1. Victim Inference
Once linked, we scan the L2 transactions for DEX swaps. A swap typically results in two ERC-20 “Transfer” events being emitted. Hence, we scan the events emitted by the identified L2 transactions and mark these as potential victims if we find two “Transfer” events emitted by the same token smart contract, where the sender and receiver of the tokens are swapped. We found a total of $1916524$ L2 transactions emitted via L1, out of which $1459570$ were observed on Arbitrum, $424212$ on Optimism, and $32742$ on zkSync. From the observed 1.9M L2 transactions, only $170674$ ($\approx$ 9%) can be considered potential victims based on our methodology to detect token swaps, where $87921$ are from Arbitrum and $82753$ from Optimism. We could not find potential victims on zkSync. This is most likely due to the fact that there was no cross-layer DEX currently deployed on zkSync at the time of our measurement. Table 9 in Appendix C provides a detailed overview of the identified DEXes and their liquidity pools on Arbitrum and Optimism. We observe that essentially all potential victim transactions originate from users using the Hop Protocol [55]. The Hop protocol allows users to send tokens from one rollup or chain to another by employing their own market makers, hence making it a perfect victim for our cross-layer sandwich attacks.
@@ 5.3.2. Transaction Inclusion Delays
Table 10 in Appendix D provides a detailed overview on the profits obtained for each attack strategy across Arbitrum and Optimism using different attacker budgets. Interestingly, not all identified potential victim transactions are profitable. These transactions are either too small to offer revenue outweighing costs, or have low slippage tolerance settings which do not provide enough room for frontrunning. In Table 6, we observe that the number of profitable transactions increases together with the attacker’s capital, suggesting that several victim transactions require large purchases by attackers in order to be profitable. Nonetheless, we can state that the maximum total profit is 1.2M USD for Arbitrum and 767K USD for Optimism. Figure 10 further highlights that Optimism is more profitable than Arbitrum for attackers that posses a capital up to 1M USD. In addition, we observe that attack strategy $S_{3}$ is always more profitable than $S_{2}$ and $S_{1}$. However, the difference diminishes as the attacker capital increases. Hence, we conclude that either strategies are profitable, but that for attackers with smaller capital (e.g., 1K USD) strategy $S_{3}$ is more profitable than for attackers with larger capital (e.g., 1M USD).
@@ 5.4. Testnet Validation
While our mainnet simulation provides insights such as number of potential victims and estimated profits, it does not validate the deployability of our attacks in practice. To that end, we leverage live test networks to demonstrate the practicality of our cross-layer sandwich attacks. Given the similarities between testnets and mainnets, we can validate the feasibility of our attacks in an ethical manner without impacting real users and requiring substantial financial resources. For our analysis, we used the Sepolia testnet alongside its corresponding L2 counterparts for Arbitrum, Optimism, and zkSync. We developed a script to automate the process of deploying our own attacker and victim accounts on each chain including our own Uniswap V2-inspired DEXes on Arbitrum, Optimism, and zkSync. The only victim was our own account even on the testnet. The script automatically performs the three attack strategies. First, a victim transaction $T_{V}$ is sent via the official L1 bridge of the rollup that is under test. The victim transaction simply performs a swap on the Uniswap-alike DEX that the script previously deployed on L2. Afterwards, the script simulates the attacker transactions $T_{A_{1}}$ and $T_{A_{2}}$ according to the selected strategy. The script either scans the mempool or the latest blocks on L1 and searches for our victim transaction $T_{V}$ as detailed in Section 5.3.1 and crafts the corresponding transactions $T_{A_{1}}$ and $T_{A_{2}}$, which essentially buy and sell the same asset as $T_{V}$ on the Uniswap-alike DEX on L2. See Table 11 in Appendix E for a detailed list of our successfully deployed attack transactions across Arbitrum, Optimism, and zkSync.
@@ 6.1. Generalization and Limitations
We performed our measurements on Arbitrum, Optimism, and zkSync, as these were the most popular rollups by market share at the time of writing [39]. However, since then other rollups such as Base [15] and Blast [12] have gained tremendously on popularity. As these rollups are based on Optimism’s OP Stack or Superchain Ecosystem [48], our scripts to detect MEV on Optimism can easily be repurposed to measure MEV extraction on any Optimism fork. While we use sandwich attacks as a case study, our proposed strategies $S_{1}$, $S_{2}$, and $S_{3}$ can be further generalized to cross-layer frontrunning attacks. We posit that our work applies any time there is a victim transaction making L1-to-L2 transactions on a blockchain with a public mempool (strategies $S_{1}$ and $S_{2}$), or when a rollup imposes a delay in including L1 transactions (strategy $S_{3}$). Sandwich attacks exploit both frontrunning and backrunning, so it is possible that we could see our strategies applied to other techniques besides sandwiching that would also yield profit when transactions are frontrun (e.g., generalized frontrunners [71, 56]). All of our proposed cross-layer sandwich attacks exploit the fact that users submit L2 transactions via L1. While this seems counter-intuitive, there exist legitimate use cases which require such behavior. For example, they are used in cross-chain swaps as provided via the Hop Protocol [55] or the Across Protocol [54]. The reason why we were only able to find victim transactions related to Hop, was due to the fact that Across leverages their own bridges and does not rely on the native bridges deployed by individual rollup technologies. However, Across’s bridge also hinges on emitting events to trigger cross-chain swaps, hence, our victim inference could be adjusted for Across as well. Strategies $S_{1}$ and $S_{2}$ depend on the fact that attackers can observe victim transactions via the mempool. Hence, any L1 blockchain that has a public mempool is susceptible to these two strategies. Finally, we did not find any victim transactions on zkSync due to the fact that during our data collection period no cross-chain exchange such as Hop was operating.